payton:

Solve this linear equation for p: 2.6(5.5p – 12.4) = 127.92 1. Distributive property: 2. Addition property of equality: 3. Division property of equality: 4. Solution: 14.3p – 32.24 = 127.92 14.3p – 32.24 + 32.24 = 127.92 + 32.24 14.3p = 160.16 StartFraction 14.3 p Over 14.3 EndFraction equals StartFraction 160.16 Over 14.3 EndFraction. p =

TheSmartOne:

One question at a time Do you know the distributive property? It's really simple a(b+c) = ab + ac So we just multiply everything inside the parenthesis by a So for example, 5(3x + 2) = (5 * 3x) + (5*2) =15x + 10

payton:

yeah i think i know the answer is it p+11.2

TheSmartOne:

So now for your question, distribute the 2.6 Tell me what you get 2.6(5.5p - 12.4) = 127.92 another version of the distributive property is a(b-c) = ab - ac So that minus sign stays

payton:

i think i know the answer to it

payton:

is it 11.2

TheSmartOne:

Bingo, you got it!

payton:

ok now i have another one

TheSmartOne:

Let's make another post for it :)

payton:

ok

TheSmartOne:

Oh I didn't realize that was all your work in the question, I thought that was a whole separate question LOL